OpenStudy (anonymous):

Approximate sqrt[3]{23.5} by means of differentials or a linear approximation. Use for the base point the closest integer that is convenient with regard to the cube root function. Warning: WeBWorK wants the approximation found using the method described above; a root read off a calculator display will be counted as wrong.

OpenStudy (anonymous):

It looks like you're asked to find an approximation for \(\sqrt[3]{23.5}\); if that's the case, I would suggest using \(a=27\) as the closest integer. So let \(f(x)=\sqrt[3]x\). The linear approximation at a point \(c=23.5\) is given by \[f(c)\approx f'(c)(c-a)+f(a)\] So, \[f(23.5)\approx f'(23.5)(23.5-27)+f(27)\] Of course, \(f(x)=\sqrt[3]x~~\Rightarrow~~f'(x)=\dfrac{3}{4}x^{4/3}\).

OpenStudy (anonymous):

so do i solve for 3/4x^3/4 to get the right answer? or the equation above that ?

OpenStudy (whpalmer4):

actually, that's not right. \(f(x) = \sqrt[3]{x} = x^{\frac{1}{3}}\)so \[f'(x) = \frac{1}{3}x^{\frac{1}{3}-1} = \frac{1}{3}x^{-\frac{2}{3}} = \frac{1}{3x^{\frac{2}{3}}}\]

OpenStudy (whpalmer4):

So \[f'(23.5) = \frac{1}{3*(23.5)^{\frac{2}{3}}}\]

OpenStudy (whpalmer4):

The result you should get if you evaluate all of that, when cubed, gives 23.3397, which is pretty close. The actual value of \(\sqrt[3]{23.5} \approx 2.86433\), for comparison purposes. Note the warning about getting a wrong answer if you are tempted to paste that in as your answer!